# Corresponding author

### v2

@@ Introduction
In this paper, we present a novel framework that utilizes a 3D engine to generate urban flood synthetic dataset, dubbed MultiFloodSynth. For fidelity, we faithfully attribute the flood hazard situation as several properties and components (e.g., layout (Shang et al. 2024), lighting, flood-level (Chaudhary et al. 2020; Wan et al. 2024), 3D object, camera view) for scene composition by exploring urban flood situation. Moreover, considering various computer vision tasks, our system includes a variety of annotation types such as normal map, instance/semantic/fine-grained segmentation map, camera 3D pose, 2D/3D bounding box, etc. Thanks to such editable attributes, our MultiFloodSynth improved the performance of object-localized flood level detection, while alleviating large dataset requirements for model training.
@@ Acknowledgments
This research was supported by Culture, Sports and Tourism R&D Program through the Korea Creative Content Agency grant funded by Ministry of Culture, Sports and Tourism in 2024 (Project Name : Developing Professionals for R&D in Contents Production Based on Generative AI and Cloud, Project Number : RS-2024-00352578).

### v3

@@ Introduction
In this paper, we present a novel framework that utilizes a 3D engine to generate urban flood synthetic dataset, dubbed MultiFloodSynth. For fidelity, we faithfully attribute the flood hazard situation as several properties and components (e.g., layout (Shang et al. 2024), lighting, flood-level (Chaudhary et al. 2020; Wan et al. 2024), 3D object, camera view) for scene composition by exploring urban flood situation. Moreover, considering various computer vision tasks, our system includes a variety of annotation types such as normal map, instance/semantic/fine-grained segmentation map, camera 3D pos, 2D/3D bounding box, etc. Thanks to such editable attributes, our MultiFloodSynth improved the performance of object-localized flood level detection, while alleviating large dataset requirements for model training.
@@ Acknowledgments
This research was supported by Culture, Sports and Tourism R&D Program through the Korea Creative Content Agency grant funded by Ministry of Culture, Sports and Tourism in 2024 (Project Name : Developing Professionals for R&D in Contents Production Based on Generative AI and Cloud, Project Number : RS-2024-00352578, Contribution Rate: 100%) and Artificial intelligence industrial convergence cluster development project funded by the Ministry of Science and ICT(MSIT, Korea) & Gwangju Metropolitan City.